A metal packs with 68% efficiency. If the edge length of the unit cell is 0.291nm, find the colume of all the atoms in this unit cell in cm^3.
1
Expert's answer
2020-09-21T06:35:13-0400

Solution.

"APF = \frac{n \times V(atom) \times 100 \%}{V(unit \ cell)}"

APF = 68 %, so it is the body-centred unit cell.

"V(atoms \ in \ unit \ cell) = \frac{APF \times V(unit \ cell)}{100 \%}"

"V(unit \ cell) = a^3"

"V(unit \ cell) = (0.291 \times 100 \times 10^{-9})^3 = 2.46 \times 10^{-23} \ cm^3"

"V(atoms \ in \ unit \ cell) = \frac{68 \times 2.46 \times 10^{-23}}{100 \%} = 1.68 \times 10^{-23} \ cm^3"

Answer:

"V(atoms \ in \ unit \ cell) = 1.68 \times 10^{-23} \ cm^3"